### v1

# LARP: Learner-Agnostic Robust Data Prefiltering

Abstract
The widespread availability of large public datasets is a key factor behind the recent successes of statistical inference and machine learning methods. However, these datasets often contain some low-quality or contaminated data, to which many learning procedures are sensitive. Therefore, the question of whether and how public datasets should be prefiltered to facilitate accurate downstream learning arises. On a technical level this requires the construction of principled data prefiltering methods which are learner-agnostic robust, in the sense of provably protecting a set of pre-specified downstream learners from corrupted data. In this work, we formalize the problem of Learner-Agnostic Robust data Prefiltering (LARP), which aims at finding prefiltering procedures that minimize a worst-case loss over a pre-specified set of learners. We first instantiate our framework in the context of scalar mean estimation with Huber estimators under the Huber data contamination model. We provide a hardness result on a specific problem instance and analyze several natural prefiltering procedures. Our theoretical results indicate that performing LARP on a heterogeneous set of learners leads to some loss in model performance compared to the alternative of prefiltering data for each learner/use-case individually. We explore the resulting utility loss and its dependence on the problem parameters via extensive experiments on real-world image and tabular data, observing statistically significant reduction in utility. Finally, we model the trade-off between the utility drop and the cost of repeated (learner-specific) prefiltering within a game-theoretic framework and showcase benefits of LARP for large datasets.
1 Introduction
The availability of large, public datasets has been crucial for the recent successes of statistical and machine learning methods. For example, public benchmarks such as ImageNET [13] and GLUE [59] have enabled the creation of numerous pre-trained and competitive models, readily available for adoption and fine-tuning by practitioners in any industry. Similarly, public healthcare datasets, e.g. the public data on the COVID-19 pandemic released by the WHO [^1] https://data.who.int/dashboards/covid19/data, can serve as a valuable reference to medical professionals for estimating statistics about common diseases and treatments.
Despite their attractiveness, public datasets often contain noisy, inaccurate or even maliciously manipulated data [8]. Since statistical methods are vulnerable to data contamination, such data issues can damage the accuracy of the downstream learning procedures applied on top of public data. For example, the Common Crawl web corpus, a widely used dataset in foundation model training, is known to contain toxic, biased, and factually incorrect content [23, 17]. Models trained on such data, including early versions of GPT-3 [7], have consequently been shown to produce toxic or biased outputs [56]. Another impact of public data inaccuracies manifested during the COVID-19 pandemic, as estimates of the disease mortality based on (hastily-collected) early public data on the virus spread were later reported to be underestimates [42].
These issues motivate the question of whether and how public datasets should be prefiltered by the data provider, so as to protect the accuracy of downstream statistical and learning procedures (hereafter referred to as learners, for brevity) applied on top of the dataset. Intuitively, principled data prefiltering can improve overall model downstream quality and also aligns with recent calls for transparency in dataset creation [22] and for data-centric strategies in developing robust and ethical AI systems [39]. However, such prefiltering poses a new technical challenge: that of designing data prefiltering algorithms that are learner-agnostic, in the sense of protecting the accuracy of a wide specification of learners simultaneously.
Contributions
In this work we study the problem of prefiltering public data with the goal of protecting a wide specification of learners, proposing a framework of Learner-Agnostic Robust data Prefiltering (LARP). Formally, the objective is to find a prefiltering, which computes a subset of an initial contaminated dataset, so as to minimize a learner-agnostic notion of risk, defined as the worst-case (maximal) loss of a set of pre-specified downstream learners trained on the prefiltered data. Differently from classic robust learning, in LARP we optimize over prefiltering procedures, instead of learners, with the best prefilter generally depending on the given specification of learners.
We first instantiate LARP in the context of scalar mean estimation with Huber contamination. We provide an example learning problem which highlights that non-trivial bounds on the learner-agnostic risk necessarily depend on the specified set of learners. Then we prove an upper bound on the learner-agnostic risk over any set of Huber estimators, achieved by a prefiltering procedure inspired by popular methods for outlier removal based on quantiles. The bound contains a term reminiscent to standard results in robust statistics, featuring the data corruption ratio and the dataset size, as well as an additional term that features the dependence on the learner set. We also observe similar behaviour of the learner-agnostic risk in empirical evaluation of three mean estimation prefiltering procedures.
Our theoretical results indicate that the learner-agnostic risk generally becomes larger with increased heterogeneity of the specified learner set, leading to a loss in utility compared to the alternative of prefiltering data for each learner/use-case individually (learner-specific prefiltering). We term the average utility reduction across the learner set as the price of learner-agnostic prefiltering and study it via extensive experiments on the Adult [3] and CIFAR-10 [36] datasets, with label noise and spurious correlations data corruptions. Finally, we model the trade-off between the price of LARP and the cost of repeated (learner-specific) prefiltering within a game-theoretic framework and showcase benefits of LARP for large datasets.
2 Related work
This work is the first, to the best of our knowledge, to present a theoretical framework for prefiltering datasets with the goal of protecting the accuracy of a pre-specified set of downstream learning algorithms. To put our work in perspective, we survey relevant works and discuss similarities and differences with our approach.
Robust statistics and learning
The problem of learning reliable estimates/models from contaminated data has long been studied in the fields of robust statistics and machine learning [29, 11]. Numerous works have studied various contamination models in the context of statistical estimation [28, 29, 33, 15, 32, 14] and supervised learning [33, 4, 16]. Some commonly studied type of data corruption in machine learning include label noise [44, 48, 26, 45, 62] and spurious correlations [24, 53, 43, 51], which we consider in our experiments.
The main difference between our setting and robust statistics/ML is that we focus on finding robust data prefiltering procedures for protecting fixed learning algorithms, while robust ML focuses on finding suitable learning algorithms directly. As we show in Sections 3 and 4, this results in a new optimization problem, which fundamentally depends on the pre-specified set of learning algorithms.
Data-centric ML
Recent advancements in data-centric machine learning emphasize the critical role of data quality in model performance. Surveys by [61] and [60] provide comprehensive overviews of methodologies aimed at improving data quality through validation, cleaning, and maintenance. One line of work studies learner-agnostic methods in the context of data valuation [31, 34]. Another approach, which is focused on mitigating label noise, is the data pruning strategy [46]. While we are also interested in prefiltering procedures that improve data quality, we study a new optimization objective for this task: that of minimizing the maximum loss of set of pre-specified learning algorithms.
Data moderation
A large body of literature studies methods for data quality control and moderation for ML data obtained via crowdsourcing [38, 2, 58, 54]. These works study specific models of data contamination, e.g. label noise, and develop techniques for data filtering with provable guarantees on the resulting data quality. Unlike these works, we only adopt limited assumptions on the type of corruption and seek to certify the quality of downstream learners directly.
Another line of work studies the impact of data moderation in the context of human learning and disinformation [25, 19, 30]. The social implications in such settings lead to many orthogonal considerations for moderators, in particular designing appropriate models of human learning, reducing polarization and diversity of the provided information. In contrast, we focus protecting the accuracy of downstream learning algorithms despite corrupted data.
Fragmented learning pipelines
Our work is motivated by the increasing availability of open-access datasets for training, which necessitates their preprocessing and preparation for public use. This is an example of the increasing fragmentation of modern learning pipelines, as multiple stakeholders become involved in one or more stages of the ML pipeline, including data gathering, preprocessing, training and deployment. In particular, several works have focused on interactions between data providers and model creators in ML. Data delegation and valuation have been studied by [9, 52, 1] with the goal of developing appropriate contract theory and fair remuneration methods for outsourcing data-related tasks. Other works consider strategic interactions between foundation model creators and entities performing fine-tuning [37]. Our work studies a concrete form of pipeline fragmentation, in which data preprocessing is performed independently of model training. We adopt the perspective of the data moderator, tasked with filtering corrupted data to ensure downstream learner robustness.
3 Framework
In this section we formalize the problem of finding learner-agnostic robust prefiltering (LARP) procedures. First, we define the environment of the problem, which consists of the contamination model and the set of downstream learners. Then we move on to present the main objective of LARP. We then discuss differences with classic robust learning and we motivate and define a notion of price of learner-agnostic prefiltering.
Contamination model
We consider the general data contamination model from classic robust learning. In the canonical setting of learning theory, one is interested in learning a property $\theta\in\Theta$ of a distribution $\mathcal{D}_{\theta}$. However, instead of i.i.d. data from the distribution $\mathcal{D}_{\theta}$, in the robust setting one assumes that the data is independently drawn from $\mathcal{D}_{\theta}$, but an arbitrary $\epsilon$-fraction of the points are contaminated according to some assumed contamination model. We refer to $\epsilon\in[0,1/2)$ as the contamination ratio and we let $\mathcal{D}^{\prime}_{\theta}$ denote the contaminated distribution.
Downstream learners
Each learner is a function $l:\cup_{n\in\mathbb{N}}^{\infty}\mathcal{X}^{n}\to\mathcal{H}$ that takes a finite sample $S$ and produces a hypothesis $\hat{h}\vcentcolon=l(S)$. For example, in the context of scalar mean estimation, we have $\mathcal{H}=\mathbb{R}$, whereas for classification we have =$\mathcal{H}=\left\{h:\mathcal{X}\to\left\{0,\ldots,c-1\right\}\right\}$. We denote the set of downstream learners as $\mathcal{L}$.
Learner-agnostic robust prefiltering
We study data prefiltering mechanisms, which filter the data before it is presented to downstream learners. Specifically, consider an $\epsilon$-contaminated dataset $S$ from the target distribution $\mathcal{D}_{\theta}$. Given this dataset, a prefiltering procedure is a function $F:\cup_{n=1}^{\infty}\mathcal{X}^{n}\to\cup_{n=1}^{\infty}\mathcal{X}^{n}$ satisfying $F(S)\subseteq S$. In other words, the dataset is mapped to one of its subsets $S^{\prime}$. The prefiltered subset $S^{\prime}$ is then presented to the set of downstream learners, where each element $l\in\mathcal{L}$ produces a hypothesis $l(S^{\prime})$. The performance of each hypothesis is then measured using a risk function $R(l(S^{\prime}),\theta)$.
Our goal is to design prefiltering procedures that protect all learners in $\mathcal{L}$. Formally, we seek to minimize a notion of learner-agnostic risk $R_{agn}(F)$ over a set of possible prefiltering procedures $\mathcal{F}$:
$\min_{F\in\mathcal{F}}R_{agn}(F),$
where
$R_{agn}(F)\vcentcolon=\max_{l\in\mathcal{L}}R(l(F(S)),\theta).$ (1)
For brevity, we denote $R_{l}(F)\vcentcolon=R\left(l\left(F\left(S\right)\right),\theta\right)$. Classic examples of risk functions include the squared loss $R(\hat{h},\theta)=(\hat{h}-\theta)^{2}$ for scalar mean estimation, or population risk $R(\hat{h},\theta)=\mathbb{P}_{(X,Y)\sim\mathcal{D}_{\theta}}(\hat{h}(X)\neq Y)$ for classification tasks. Note that the learner-agnostic risk presented in Equation 1 is a random variable since it depends on the random sample $S$. In this work, we focus on high-probability bounds for $R_{agn}$ in our theoretical results.
Analyzing the learner-agnostic risk
Although our setup is concerned with producing accurate hypotheses, similarly to classical robust learning, there are two important conceptual differences. We can formally define robust learning as finding $\operatorname*{argmin}_{l}R\left(l(S),\theta\right)$ over all possible learners $l:\cup^{\infty}_{n=1}\mathcal{X}^{n}\to\mathcal{H}$. The main difference with our framework is that in the LARP problem the learner set $\mathcal{L}$ is fixed, and instead minimization happens over the prefiltering procedure $F$. Thus, the optimal procedure may vary significantly according to the properties of the downstream learners. This can be further highlighted by rewriting Equation 1 as
$R_{agn}(F)$ :$\vcentcolon=\max_{l\in\mathcal{L}}R_{l}(F)=\min_{l\in\mathcal{L}}R_{l}(F)+\left(\max_{l\in\mathcal{L}}R_{l}(F)-\min_{l\in\mathcal{L}}R_{l}(F)\right).$ (2)
The first term $\min_{l\in\mathcal{L}}R_{l}(F)$ denotes the best possible error that can be achieved by any learner in the set. Under the assumption that at least one learner in $\mathcal{L}$ is reasonable, this resembles the target objective of classic robust learning. The second term encapsulates the hardness that arises from the heterogeneity between the learners. In the general case, this term is positive and induces additional losses not present in standard robust frameworks. In Section 4 we study the objective $R_{agn}(F)$ in further detail for the case of scalar mean estimation.
Price of learner-agnostic prefiltering
The second term in Equation 2 indicates that applying LARP for larger learner sets may lead to worse downstream performance. This is intuitive, as prefiltering a public dataset with the goal of protecting the performance of multiple learning procedures is intuitively harder than prefiltering the dataset for a specific learner only (e.g. prefiltering performed by anyone using the dataset for their own use-case).
To quantify the loss in utility resulting from such accuracy drop, we introduce a measure of “price of learner-agnostic prefiltering”. Formally, we call an instance of LARP learner-specific prefiltering for $l$ if =$\mathcal{L}=\left\{l\right\}$. We want to compare the across-learners performance of a prefiltering procedure $F\in\mathcal{F}$ selected by LARP on a learner set $\mathcal{L}$, to the performance of the learner-specific optimal prefiltering procedures $F^{\ast}_{l}=\operatorname*{argmin}_{F\in\mathcal{F}}R_{l}\left(F\right)$ for each $l\in\mathcal{L}$. Note that $R_{l}(F)\geq R_{l}(F_{l}^{\ast}),\forall F\in\mathcal{F},l\in\mathcal{L}$, which leads to each learner losing utility which we describe using a function $\mathcal{U}_{red}:\mathbb{R}^{2}\to\mathbb{R}$  [^2] For fixed $R_{1},R_{2}$, the value $\mathcal{U}_{red}(R_{1},R_{2})$ is the reduction in learner $l$ ’s utility if their risk is $R_{1}$ instead of $R_{2}$.. Then we define the price of learner-agnostic prefiltering of a prefiltering procedure $F$ as
$P\left(F\right)\vcentcolon=\frac{1}{\lvert\mathcal{L}\rvert}\sum_{l\in\mathcal{L}}\mathcal{U}_{red}\left(R_{l}(F),R_{l}(F_{l}^{\ast})\right).$ (3)
Despite the loss in utility, LARP can still be beneficial as it is intuitively “cheaper” to prefilter a dataset once, than for each learner individually. Therefore, in Section 5 we measure the price of learner-agnostic prefiltering on real data and study incentives for performing LARP over learner-specific prefiltering.
4 LARP for scalar mean estimation
In this section we present a theoretical analysis of LARP for the problem of scalar mean estimation. We begin by discussing a particular instance of the framework that highlights the additional loss incurred by LARP compared to classic robust learning, drawing parallels to the decomposition presented in Equation 2. Afterwards, we present guarantees for a prefiltering procedure inspired by popular rules for outlier-removal based on quantiles. Finally, we empirically study three prefiltering mechanisms for the task of Gaussian mean estimation.
4.1 Setup
In this section we focus on scalar mean estimation within the Huber contamination model.
Huber contamination
We assume the standard Huber contamination [28], in which the $\epsilon$-contaminated sample $S$ is drawn i.i.d. from a distribution $\mathcal{D}^{\prime}_{\theta}=(1-\epsilon)\mathcal{D}_{\theta}+\epsilon Q,$ where $Q\in\mathfrak{Q}$ is a noise distribution. Note that since $Q$ is unknown in practice any upper bound on $R_{agn}$ should hold for all possible $Q\in\mathfrak{Q}$.
Huber estimators
The set of learners $\mathcal{L}$ is a set of Huber estimators [28, 57, 50], which are parametrized by a scalar $\delta>0$. They are defined as $\hat{\theta}_{\delta}(S)\vcentcolon=\operatorname*{argmin}_{\hat{\theta}\in\mathbb{R}}\sum_{x\in S}H_{\delta}(x-\hat{\theta})$, where the Huber loss $H_{\delta}$ is defined as
$H_{\delta}(x)\vcentcolon=\begin{cases}\frac{1}{2}x^{2}&\text{for $\lvert x\rvert\leq\delta$}\\ \delta\left(\lvert x\rvert-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$
In the cases where $\delta\to 0$ and $\delta\to\infty$, the Huber estimator converges to the sample median and the sample mean respectively. The intermediate values of $\delta$ present the tradeoff between robustness and sample efficiency. We denote the set of the parameters of the given Huber learners as $\Delta\vcentcolon=\left\{\delta:H_{\delta}\in\mathcal{L}\right\}$.
Finally, we define a risk function that measures the square distance between the estimate and the true mean: $R\left(l(S^{\prime}),\theta\right)\vcentcolon=(l(S^{\prime})-\theta)^{2}$.
4.2 Hardness of learning with high heterogeneity
The heterogeneous behaviour of the downstream learners can lead to high losses, no matter the prefiltering. To show this we provide a lower bound on the learner-agnostic risk, by considering a specific example of a target dataset, noise, and learner set.
Lemma 4.1.
There is an instance of LARP with specified $\mathcal{D}_{\theta}=Ber(1/2)$ and fixed $\mathcal{L}$ such that: 1) there exists a prefiltering with $\min_{l\in\mathcal{L}}R_{l}=\mathcal{O}(\epsilon^{2})$, but 2) for all prefiltering procedures, $R_{agn}=\Omega(1)$.
We refer to Section A.1 for the construction, proof and a more thorough discussion. This lower bound arises due to the presence of learners that are inefficient for the instance. Therefore, while the first term in Equation 2 can be made small, this will cause the second term to be large. This is a strong indicator that any non-trivial bound on the learner-agnostic risk needs to be a function of the learner set $\mathcal{L}$.
4.3 Prefiltering mechanism based on quantiles
We now prove feasibility of LARP in the context of a scalar mean estimation for a Gaussian distribution $\mathcal{N}(\theta,\sigma^{2})$, by providing a prefiltering procedure with provable upper bounds on its learner-agnostic risk. This procedure is inspired by a popular informal rule for marking the tails of a sample as outliers [41]. This can be adapted to define the following outlyingness measure:
$Q\left(x,\left\{X_{1},\ldots,X_{n}\right\}\right)=\frac{1}{n}\left\lvert\min\left\{i:X_{(i)}\geq x\right\}-\frac{n}{2}\right\rvert.$
Based on that, we can define a quantile prefiltering procedure as
$F^{q}_{p}(S)\vcentcolon=\left\{X\in S:Q(X,S)<p\right\}.$
Sensible values for the hyperparameter $p$ are in the range $\left(0,1/2\right)$. The following result provides guarantees for the learner-agnostic risk of $F^{q}_{p}$.
Theorem 4.2.
Assume that the target distribution is $\mathcal{D}_{\theta}=\mathcal{N}(\theta,\sigma^{2})$, and that $\epsilon<2/7$. Let $F_{p}$ be the quantile prefiltering procedure with any $p\in(0,1/2)$. Then, if $n\geq\Omega(\log(1/\delta_{0}))$, with probability $1-\delta_{0}$ the downstream Huber learners with parameter set $\Delta$ produce mean estimates $\hat{\theta}_{\delta}$ such that
$\max_{\delta\in\Delta}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}\leq\mathcal{O}\left(\left(\epsilon^{2}+\log(1/\delta_{0})/n\right)\sigma^{2}+\max_{\delta\in\Delta}\delta^{2}\right).$ (4)
The first term in Equation 4 is reminiscent of upper bounds in classic robust mean estimation (see e.g. [14]) and increases with the corruption rate $\epsilon$ and decreases with the sample size. The second term depends solely on the learner set and is always nonnegative.
We note that small values of $\delta\approx 0$ yield estimators close to the sample median, which is known to enjoy guarantees similar to the first term in Equation 4 [14]. Therefore, if all estimators in $\mathcal{L}$ use a small value of $\delta$ (i.e. all estimators are “good” for the considered learning problem), the bound is comparable to those in standard robust mean estimation. However, if some estimators use a large $\delta$, they are suboptimal for the problem and the bound is also larger as $\max_{\delta\in\Delta}\delta^{2}$ is large. This aligns with the arguments presented in Section 4.2, since any non-trivial bound on $R_{agn}$ must also take into account the set of downstream learners.
4.4 Empirical analysis of the Gaussian mean estimation setup
In this subsection we empirically explore the dependence of $R_{agn}$ on $\epsilon$ and $\Delta$ with $\mathcal{D}_{\theta}=\mathcal{N}(0,1)$, by analyzing three different prefiltering procedures. The first one is $F_{p}^{q}$ as defined in Section 4.3, the other two are analogously defined using z-score [41] and Stahel-Donoho outlyingness (SDO) [55, 18, 63]. Each procedure is being optimized over their own hyperparameter that controls the amount of data being removed. Full details of the experiments are described in Section B.1.
In Figure 1(a) we notice that the empirical trend exhibits superlinear growth, in alignment with the $\epsilon^{2}$ term in Theorem 4.2. This is also consistent with lower bounds that we find in [14]. We also observe that $F^{z}$ and $F^{sdo}$ outperform $F^{q}$ in this setup. In any case, our results show that learning is viable under the assumption of reasonable downstream learners. In Figure 1(b) we measure the second term in Equation 2 on a set =$\mathcal{L}=\left\{\delta_{1},\delta_{2}\right\}$ where $\delta_{1}=0.01$ and $\delta_{2}$ varies along the x-axis, thus creating varying levels of heterogeneity between the learners. We see that increasing learner heterogeneity leads to larger additional loss incurred by the models. Moreover, the behaviour of this term is different for the different prefiltering procedures, suggesting that the optimal choice might depend on the learner set. This aligns with the analysis presented in Section 4.2, where we argued that the design of procedures for LARP should depend on the downstream learner set. In all cases, we conclude that the effect of the learner set is statistically significant and the guarantees for downstream learners are inherently worse due to learner heterogeneity.
(a) Learner-agnostic risk as a function of $\epsilon$.
(b) Gap between best-case and worst-case risk.
Figure 1: Experiments for Gaussian mean estimation.
5 Price of learner-agnostic prefiltering
Previous results show that performing LARP on a non-trivial set of learners leads to a reduction of downstream performance compared to learner-specific instances. The average reduction over the learner set can be measured through the notion of price of learner-agnostic prefiltering as defined in Section 3. In this section we measure the price of learner-agnostic prefiltering for different real-world tasks, with various types of data contamination. We conduct ablation studies to investigate the dependence on contamination ratio, learner heterogeneity and dataset size. Finally, we present a theoretical argument for the benefits of learner-agnostic prefiltering for large datasets.
5.1 Price of learner-agnostic prefiltering on real-world data
In this section, we quantify the price of learner-agnostic prefiltering on datasets with inherent noise. Specifically, we conduct experiments in both the tabular and image classification settings, using two realistic noise models: i) Uniform Label Noise, which mimics scenarios where annotation is unreliable; and ii) Shortcuts [24], which simulates scenarios where training data contains spurious correlation features irrelevant to the labels, e.g. textual information embedded in the CT image patches. In all cases, the price $P$ is computed using the utility function $\mathcal{U}_{red}(R_{1},R_{2})\vcentcolon=(R_{1}-R_{2})/R_{2}$, that measures the relative increase in classification risk when learner-agnostic prefiltering is used.
Datasets
For image experiments we use CIFAR-10 [36], splitting its train set into train (80%) and validation (20%) sets and using the original test set. Tabular experiments utilize Adult [3], randomly split into train (64%), validation (16%), and test (20%) sets. The validation sets are utilized to select the hyperparameters (Adult), and perform early-stopping (CIFAR-10). In all experiments, we apply the noise to the train and validation sets, but evaluate on clean data to measure learner generalization.
Label noise
In label noise experiments, we use the model accuracy as our risk metric $R$, $\epsilon=30\%$ uniform label noise on all classes, and prefilter up to $p\%$ of the train and validation sets. Motivated by the practical observation that complex models tend to learn useful features first before overfitting to noisy data, the prefiltered samples were selected by maximizing the loss of chosen overparametrized models (ResNet-9 [27] on CIFAR-10, and 2 layer fully connected NN on Adult) trained on the combined train and evaluation sets with early-stopping. We found this procedure effective in practice, resulting in $>75\%$ of corrupted samples being filtered at the default value of $p=25\%$.
Shortcuts
For tabular data, we inject noise by maximizing the integer feature “education-num” in $85\%$ of the data for the class corresponding to the label “$>$50K”, leading to $\epsilon\approx 21\%$ contamination rate due to class imbalance. For image data, we set the color of a $3\times 3$ patch in the lower right corner of $85\%$ of the data in classes 0,2,4 according to their label, resulting in $\epsilon=25.5\%$ contamination rate. Based on the observation that models learn shortcut data quickly, in both cases we prefilter the top-$p$ percent of lowest-loss datapoints based on CNN (CIFAR-10) and Random Forests (Adult) models from the respective learner sets, but trained on the combined train and validation sets. We verify that this results in $>85\%$ of prefiltered shortcut data in practice at the default $p=25\%$. We use the macro-F1 score [40] as our risk metric $R$ due to the class imbalance of the prefiltered datasets.
Learner sets
We evaluate LARP on sets of models with different noise robustness levels, which can intuitively be seen as the worst-case heterogeneity, as models with high robustness will achieve optimal risk under little-to-no prefiltering, while models with low robustness will achieve optimal risk under heavy prefiltering. As we train very large number of models, for CIFAR-10 we opt to use convolutional networks (CNNs) with an architecture (See Section B.2) optimized for fast training that still maintains good accuracy. As we observe that L2 regularization increases the robustness to noisy labels in practice, but is also susceptible to underfitting, in the label noise setting, we generate our set of models $\mathcal{L}$ by varying the CNNs’ L2 regularization parameter in the range $[3e{-3},2e{-2}]$. For shortcut experiments we regularize the effect of the shortcuts on the training explicitly by reducing the gradient contribution of the shortcut patch pixels by a factor in the range $[1e{-6},1]$. For Adult experiments, we opt to use learners based on diverse set of machine learning model types — XGBoost [10], AdaBoost [20], LogitBoost [21], Bagging [5], RandomForest [6], SVM [12], as well as neural networks with two hidden layers and ReLU activations; in order to show that our conclusions generalize beyond families of models differing only by a single continuous parameter. For each model type we select two sets of hyperparameters — robust and brittle, resulting in 14 total models in $\mathcal{L}$. The robust ones were selected by minimizing the noisy validation data risk on 20 random hyperparameter configurations, while the brittle models were optimized to increase the difference between the risks on the prefiltered validation sets of two versions of the model – one trained on prefiltered train data, and one trained on noisy train data. Our brittle models intuitively represent learners that chose their parameters for maximal accuracy, disregarding robustness completely.
We train all neural networks using PyTorch [47] with Adam optimizer [35] and the remaining models are trained using Scikit-learn [49]. In all shortcut experiments learners reweight their losses to account for class imbalance of the prefiltered sets. Full experimental details can be found in Section B.2.
Results
In this paragraph, we show how the price of LARP, $P$, is affected by the different factors in the equation of the learner-agnostic risk $R_{agn}$ in Theorem 4.2. In particular, in Figure 2 we show the effects of the contamination ratio $\epsilon$, the dataset size $n$, and the learner heterogeneity, expressed in terms of the average price $P$ on all subsets of size $k$ of the learner set $\mathcal{L}$.
Crucially, we see a statistically significant price of learner-agnostic prefiltering $P$, confirming practically our theoretical results in Theorem 4.2. Moreover, our results are consistent across setups, suggesting that the price of learner-agnostic prefiltering can be incurred for different data modalities, prefiltering procedures $\mathcal{F}$, and learner sets $\mathcal{L}$.
We observe that heterogeneity has the most pronounced effect on the price of LARP, confirming our theoretical finding that LARP in the presence of heterogeneous learner set $\mathcal{L}$ is a fundamentally harder problem compared to classical robust learning. In Appendix C, we experiment with a different notion of learner heterogeneity — learner diameter that measures the range of regularization parameters used by the set of learners in $\mathcal{L}$, again confirming noticeable heterogeneity impact on $P$.
Another important factor for the price $P$ is the contamination ratio $\epsilon$. In our experiments, we observe that in general $P$ increases together with $\epsilon$, especially in the shortcut setting. We postulate this is because when more and more shortcuts are not prefiltered, the non-robust models become more and more susceptible to them, thus driving the overall price of learner-agnostic prefiltering up. We see less pronounced effect of $\epsilon$ in the label noise settings where too large noise even seems to decrease $P$. This suggests, that the relationship to $\epsilon$ is complex and likely dependent on the particular prefiltering procedure, explaining why our theory does not account for it. Theoretically exploring this relationship further is an exciting future work item.
Finally, we explore the relations of the price of LARP to the datasize $n$. For the most part, the effect appears to be minimal, and in many cases, the price of LARP even decreases with increasing $n$. In Appendix D we present results for $P$ in the context of mean estimation which are in agreement with this relationship. This motivates us to further explore the practical benefits of LARP on large datasets in Section 5.2.
Adult (Label) Adult (Shortcut) CIFAR-10 (Label) CIFAR-10 (Shortcut)
Heterogeneity
Refer to caption Refer to caption Refer to caption Refer to caption
Contam. Rate
Refer to caption Refer to caption Refer to caption Refer to caption
Dataset Size
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Price of learner-agnostic prefiltering for Adult (left two columns) and CIFAR-10 (right two columns). Each column corresponds to a dataset and corruption type (label or shortcut). From top to bottom: heterogeneity, contamination ratio, and dataset size.
5.2 Benefits from learner-agnostic prefiltering
In our theory and experiments, we have observed that LARP leads to a loss in utility resulting from an average accuracy drop of the learners, compared to learner-specific prefiltering. Nevertheless, prefiltering large datasets for each learner individually (i.e. performing learner-specific prefiltering) is intuitively impractical, as the very process of prefiltering can be costly.
In this section we study a game-theoretic model that captures this trade-off between the price of LARP and the cost of repeated prefiltering. In our setting, each learner can either choose learner-specific prefiltering and pay the full cost of preprocessing the dataset, or take part in learner-agnostic prefiltering and split said cost with other learners. We provide a theoretical result which provides a sufficient condition under which all learners are incentivized to participate in LARP.
Definition
Assume that we conduct the prefiltering setup as presented in Section 3. We consider a game in which each of the downstream learners $l\in\mathcal{L}$ is a player trying to maximize their utility. Each learner $l$ has utility $U^{l}$ that is a function of their risk $R_{l}(F)$, as well as the cost they pay for the prefiltering procedure. We model the total cost of prefiltering a dataset of size $n$ as $Cn^{\alpha}$ for some constants $C>0,\alpha\geq 1$.
Each learner selects an action $a_{l}\in\left\{0,1\right\}$ indicating whether they want to take part in a learner-specific or learner-agnostic procedure respectively. If learner $l$ plays $a_{l}=0$, they receive individual prefiltering, which costs $Cn^{\alpha}$, and receive risk $R_{l}(F^{\ast}_{l})$. Let $U^{l}_{spec}$ denote their final utility in this scenario.
On the other hand, if learner $l$ plays $a_{l}=1$, then they participate in learner-agnostic prefiltering with other players. Then, all learners can distribute the cost $Cn^{\alpha}$ of prefiltering among them according to a vector $(p_{l})_{l\in\mathcal{L}}$, i.e. $\sum_{l\in\mathcal{L}}p_{l}=Cn^{\alpha}$. In this case, learner $l$ loses $\mathcal{U}_{red}\left(R_{l}(F),R_{l}(F_{l}^{\ast})\right)$ utility due to reduced performance, but they also reduce their cost of prefiltering from $Cn^{\alpha}$ to $p_{l}$ Then the utility for player $l$ becomes $U^{l}_{agn}=U^{l}_{spec}-\mathcal{U}_{red}\left(R_{l}(F),R_{l}(F_{l}^{\ast})\right)+Cn^{\alpha}-p_{l}$. These utilities are tightly connected to the price of learner-agnostic prefiltering defined in Equation 3. The following result provides a sufficient condition for the benefit of LARP over learner-specific prefiltering.
Lemma 5.1.
Assume that in the aforementioned setup, the dataset size $n$ satisfies
$n>\left[\frac{\lvert\mathcal{L}\rvert}{C(\lvert\mathcal{L}\rvert-1)}P(F)\right]^{1/\alpha}.$
Then, there is a payment scheme $(p_{l})_{l\in\mathcal{L}}$ such that $U^{l}_{agn}\geq U^{l}_{spec}\text{ for all }l\in\mathcal{L}$. In other words, no learner is incentivized to opt out of the learner-agnostic prefiltering scheme.
Furthermore, this condition holds for sufficiently large $n$ as long as $P(F)$ grows as $o(n^{\alpha})$. Harking back to Figure 2, we observe indications of sublinear growth across all setups. This gives us an empirical confirmation of the overall benefit from LARP.
6 Conclusion
In this work, we studied the problem of prefiltering public data with learner-agnostic guarantees. We presented a formal framework for the problem and highlighted differences to classic robust learning. We presented a specific instance of the framework that showcases the need for guarantees that depend on the learner set. We then proved upper bounds on the achievable maximum risk for the task of mean estimation for scalar Gaussian random variables. We also conducted an empirical analysis on real-world tabular and image data in order to measure the effect of learner heterogeneity on downstream performance. Finally, we argued about the benefits of LARP on large datasets.
Limitations and future work
We see our work as an initial step towards understanding how prefiltering datasets can provide reliability guarantees for inference and learning procedures applied on top of them. We consider generalizations of the results presented in Section 4 to regression and classification tasks, as well as the design of provably optimal prefiltering procedures, to be an exciting direction for future work.
Acknowledgements
This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure). This project was supported with computational resources provided by Google Cloud Platform (GCP).
Supplementary Material
The supplementary material is structured as follows.
  • Appendix A contains proofs of theoretical results.
  • Appendix B contains further experimental details.
  • Appendix C contains additional experiments on real-world data.
  • Appendix D contains theoretical results for price of learner-agnostic prefiltering in the context of mean estimation.
  • Appendix E contains miscellaneous information.
Appendix A Proofs of results presented in Section 4
A.1 Proof and discussion of Lemma 4.1
We first restate and prove Lemma 4.1.
Lemma 4.1.
There is an instance of LARP with specified $\mathcal{D}_{\theta}=Ber(1/2)$ and fixed $\mathcal{L}$ such that: 1) there exists a prefiltering with $\min_{l\in\mathcal{L}}R_{l}=\mathcal{O}(\epsilon^{2})$, but 2) for all prefiltering procedures, $R_{agn}=\Omega(1)$.
Proof.
Recall that the prefiltering procedure aims to minimize the loss presented in Equation 1. Note that for a finite sample size the prefiltering procedure induces a distribution on the filtered dataset. By letting the sample size size go to infinity, this will lead to a limit distribution, which we will denote as $F(P^{\prime}_{\theta})$. We will refer to this limit regime as “infinite sample size”. We show an example that has high risk even in the infinite sample size regime.
Assume that the target distribution is $Ber(\theta)$, where either $\theta=(1-\epsilon)/2$ or $\theta=(1+\epsilon)/2$, where $\epsilon$ is the contamination ratio. Furthermore, assume that the contamination is such that the contaminated distribution is $P^{\prime}_{\theta}=Ber(1/2)$. Finally, assume that the downstream learners are Huber estimators with parameters $\delta=0,\delta=1/4,\delta=2$. The first one corresponds to mean absolute loss minimizer, the last one corresponds to the mean [^3] Due to the support of the particular distributions, each datapoint will be considered with square loss when $\delta=2$., and the middle one corresponds to some intermediate estimator. Let us denote the corresponding estimates by $\hat{\theta}_{0},\hat{\theta}_{1/4}$ and $\hat{\theta}_{2}$ respectively. This instance of the framework is presented in LABEL:fig:lower_bound. Then, any prefiltering procedure $F$ receives the contaminated distribution $P^{\prime}_{\theta}$ and maps it to a distribution $F(P^{\prime}_{\theta})$. Moreover, since the procedure is only allowed to filter points, the probability mass function $p^{\prime}$ of $P^{\prime}_{\theta}$ must be absolutely continuous w.r.t. the probability mass function $p$ of $P_{\theta}$. In particular, this means that $P^{\prime}_{\theta}$ is also Bernoulli distributed with some parameter $p_{1}$. We can describe each possible $F$ with the parameter $p_{1}\in[0,1]$ in the prefiltered distribution. Then, we can explicitly calculate the values of the estimators as a function of $p_{1}$. In particular, if $p_{1}>1/2$, then
$\hat{\theta}_{0}=1,\hat{\theta}_{1/4}=1-\frac{1-p_{1}}{4p_{1}},\hat{\theta}_{2}=p_{1}.$
In the other case, $p_{1}\leq 1/2$, then
$\hat{\theta}_{0}=0,\hat{\theta}_{1/4}=\frac{p_{1}}{4(1-p_{1})},\hat{\theta}_{2}=p_{1}.$
We begin by noting that $\min_{l}R_{l}(F)$ can be shown to be $\mathcal{O}(\epsilon^{2})$, by selecting a prefiltering mechanism that balances out the mass at 00 and $1$, i.e. returning the distribution $Ber(1/2)$, and then consider the sample mean estimator $\theta_{2}$.
Nevertheless, the learner-agnostic risk $R_{agn}$ that we aim to minimize satisfies
$R_{agn}(p_{1})=\max_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}\geq\Omega(1)$
for all $p_{1}\in[0,1]$. ∎
The reason for this difference is the existence of estimators which can be considered to be suboptimal for the given task and distribution, and also the difference between the estimates that are given here. Returning to the decomposition presented in Equation 2, we can see that each of the second and third terms can be made small at the expense of the other. For example, if we pick $p_{1}=1/2$, then we have
$\min_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}\leq\mathcal{O}(\epsilon^{2}).$
Nevertheless, the second term, corresponding to learner heterogeneity, satisfies
$\max_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}-\min_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}\geq\Omega(1).$
On the other hand, the last term in Equation 2 is not bounded from below by any positive constant. Indeed, it can always be reduced to zero by using a prefiltering procedure that collapse the distribution to some delta function. In this case all reasonable estimates will return that constant data point as an estimate that passes through $F$. Of course, in the general case this is not useful since a constant prefiltered distribution leads to huge loss of information about the initial $\theta$. Harking back to our example, if we select $p_{1}=0$ or $p_{1}=1$, then we will achieve zero learner heterogeneity, i.e.
$\max_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}-\min_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}=0.$
Nevertheless, this results in the fact that all estimates are bad in the sense that
$\min_{\delta\in\left\{0,1/4,2\right\}}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}=\Omega(1).$
The conclusion that we take from this example is that all meaningful bounds of the learner-agnostic framework must take into account not only the robustness properties of the statistical task, but also the downstream learners, since if they are sufficiently inefficient or non-robust, learning is simply not possible.
A.2 Proof of Theorem 4.2
Theorem 4.2.
Assume that the target distribution is $\mathcal{D}_{\theta}=\mathcal{N}(\theta,\sigma^{2})$, and that $\epsilon<2/7$. Let $F_{p}$ be the quantile prefiltering procedure with any $p\in(0,1/2)$. Then, if $n\geq\Omega(\log(1/\delta_{0}))$, with probability $1-\delta_{0}$ the downstream Huber learners with parameter set $\Delta$ produce mean estimates $\hat{\theta}_{\delta}$ such that
$\max_{\delta\in\Delta}\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}\leq\mathcal{O}\left(\left(\epsilon^{2}+\log(1/\delta_{0})/n\right)\sigma^{2}+\max_{\delta\in\Delta}\delta^{2}\right).$ (5)
Proof.
Corollary 1.15 of [14] states that if $\delta_{0}$ is at least $e^{-an}$ for sufficiently small $a$, then with probability $1-\delta_{0}$ we have
$\lvert m_{med}-\theta\rvert\leq\mathcal{O}\left(\epsilon+\sqrt{\log(1/\delta_{0})/n}\right)\sigma,$
where $m_{med}$ is the median of the contaminated distribution. Moreover, the quantile mechanism always preserves the median, hence the median $m^{\prime}_{med}$ of the prefiltered sample retains the above property.
We now focus on providing an upper bound on the distance between a Huber estimator $\hat{\theta}_{\delta}$ and the sample median $m^{\prime}_{med}$. First, we note that the Huber estimator minimizes the loss $\sum_{x\in S^{\prime}}H_{\delta}(x-\hat{\theta})$. This is equivalent to solving the equation $\sum_{x\in S^{\prime}}H^{\prime}_{\delta}(x-\hat{\theta})=0$, where
$H^{\prime}(x)=\begin{cases}x&\text{for $\lvert x\rvert\leq\delta$}\\ \delta sign(x)&\text{otherwise.}\end{cases},$
where $sign$ denotes the sign function. Note in particular that $H^{\prime}_{\delta}(x)\in[-\delta,\delta]$ for all $x\in\mathbb{R}$. Hence, if $\hat{\theta}_{\delta}-m^{\prime}_{med}>\delta$, we are bound to have positive value for $\sum_{x\in S^{\prime}}H^{\prime}_{\delta}(x-\hat{\theta}_{\delta})$. Similar reasoning goes for the case where $\hat{\theta}_{\delta}-m^{\prime}_{med}<-\delta$. Hence, for any Huber learner $\hat{\theta}_{\delta}$ with parameter $\delta$ trained on the prefiltered sample $F(S^{\prime})$ we have the bound
$\lvert\hat{\theta}_{\delta}-m^{\prime}_{med}\rvert\leq\delta.$
Combining the above bounds we get
$\lvert\hat{\theta}_{\delta}-\theta\rvert^{2}\leq 2\lvert\hat{\theta}_{\delta}-m^{\prime}_{med}\rvert^{2}+2\lvert m^{\prime}_{med}-\theta\rvert^{2}\leq\mathcal{O}\left(\delta^{2}+\left(\epsilon^{2}+\log(1/\delta_{0})/n\right)\sigma^{2}\right).$
This gives us the final upper bound on the worst-case estimator
$\max_{\delta\in\Delta}\lvert\hat{\theta}_{\delta}-\theta_{\mathcal{D}}\rvert^{2}\leq\mathcal{O}\left(\max_{\delta\in\Delta}\delta^{2}+\left(\epsilon^{2}+\log(1/\delta_{0})/n\right)\sigma^{2}\right).$
A.3 Proof of Lemma 5.1
Lemma 5.1.
Assume that in the aforementioned setup, the dataset size $n$ satisfies
$n>\left[\frac{\lvert\mathcal{L}\rvert}{C(\lvert\mathcal{L}\rvert-1)}P(F)\right]^{1/\alpha}.$
Then, there is a payment scheme $(p_{l})_{l\in\mathcal{L}}$ such that $U^{l}_{agn}\geq U^{l}_{spec}\text{ for all }l\in\mathcal{L}$. In other words, no learner is incentivized to opt out of the learner-agnostic prefiltering scheme.
Proof.
Suppose that the condition
$n>\left[\frac{\lvert\mathcal{L}\rvert}{C(\lvert\mathcal{L}\rvert-1)}P(F)\right]^{1/\alpha}$
is satisfied. Let us pick a cost distribution $(p_{l})_{l\in\mathcal{L}}$ such that
$p_{l}\propto Cn^{\alpha}-\mathcal{U}_{red}(R_{l}(F),R_{l}(F_{l}^{\ast}))$
and $\sum_{l\in\mathcal{L}}p_{l}=Cn^{\alpha}.$ As noted in Section 5.2, we have the difference in utilities between learner-agnostic and learner-specific prefiltering for learner $l$ is
$U^{l}_{agn}-U^{l}_{spec}=Cn^{\alpha}-p_{l}-\mathcal{U}_{red}(R_{l}(F),R_{l}(F_{l}^{\ast})).$
On the other hand, we have for all $l\in\mathcal{L}$ the bound
$\sum_{l\in\mathcal{L}}\left[Cn^{\alpha}-\mathcal{U}_{red}(R_{l}(F),R_{l}(F_{l}^{\ast}))\right]$ $=\lvert\mathcal{L}\rvert Cn^{\alpha}-\sum_{l\in\mathcal{L}}\mathcal{U}_{red}(R_{l}(F),R_{l}(F_{l}^{\ast}))$
$=\lvert\mathcal{L}\rvert Cn^{\alpha}-\lvert\mathcal{L}\rvert P(F)$
$\geq\lvert\mathcal{L}\rvert Cn^{\alpha}-C\left(\lvert\mathcal{L}\rvert-1\right)n^{\alpha}$
$=Cn^{\alpha}.$
Hence, we have for all $l\in\mathcal{L}$ the inequality
$p_{l}=\frac{Cn^{\alpha}-\mathcal{U}_{red}(R_{l}(F),R_{l}(F_{l}^{\ast}))}{\sum_{l^{\prime}\in\mathcal{L}}\left[Cn^{\alpha}-\mathcal{U}_{red}(R_{l^{\prime}}(F),R_{l^{\prime}}(F_{l^{\prime}}^{\ast}))\right]}Cn^{\alpha}\leq Cn^{\alpha}-\mathcal{U}_{red}(R_{l}(F),R_{l}(F_{l}^{\ast})).$
This implies that $U^{l}_{agn}\geq U^{l}_{spec}$ for all $l\in\mathcal{L}$, as desired.
Appendix B Further experimental details
Section B.1 contains full details of the experiments presented in Section 4.4. Section B.2 contains additional details of the experiments presented in Section 5.1.
B.1 Additional details for experiments on mean estimation
We conduct an experiment on an instance of LARP in the context of mean estimation with $\mathcal{D}_{\theta}=\mathcal{N}(0,1)$ and hence $\theta=0$. The sample size we set is $n=10001$. We simulate Huber contamination by using a set of noise distributions $\mathfrak{Q}=\left\{\mathcal{N}(m,1):m\in\mathbb{R}\right\}$. We can use symmetry arguments to restrict the space to $m\geq 0$, and we empirically observe that considering $50$ equidistant values in $\left[0,10\right]$ for $m$ recovers the worst-case contamination in all experiments.
The first prefiltering is defined as in Section 4.3. The second prefiltering that we study is based on the famous “three-sigma edit” rule . Intuitively speaking, given a sample $x_{1},\ldots,x_{n}$, we can measure the “outlyingness” of a single observation as $t_{i}=(x_{i}-\overline{x})/s$ and then removing all points for which $\lvert t_{i}\rvert>3$. This measure is also known as the z-score of the observation. Motivated by this, we study the sample z-score function as
$Z\left(x,\{X_{1},\ldots,X_{n}\}\right)\vcentcolon=\left\lvert\frac{x-M(X_{1},\ldots,X_{n})}{SD(X_{1},\ldots,X_{n})}\right\rvert$
where the sample mean $M(X_{1},\ldots,X_{n})$ and the sample standard deviation $SD(X_{1},\ldots,X_{n})$ are defined as
$M(X_{1},\ldots,X_{n})$ $=\frac{1}{n}\sum_{k=1}^{n}X_{n}$
$SD(X_{1},\ldots,X_{n})$ $=\sqrt{\frac{1}{n}\sum_{k=1}^{n}\left(X_{i}-M(X_{1},\ldots,X_{n})\right)^{2}}.$
Then, we can define the z-score prefiltering mechanism as
$F^{z}_{l}\left(S\right)\vcentcolon=\left\{X\in S:Z\left(X,S\right)<l\right\}.$
Note that this mechanism has a hyperparameter $l$ that can be selected in order to apply a “$l$-sigma edit” rule on a dataset.
The third prefiltering mechanism is based on Stahel-Donoho outlyingness (SDO), which can be defined as
$SDO\left(x,\left\{X_{1},\ldots,X_{n}\right\}\right)\vcentcolon=\frac{\lvert x-Med(X_{1},\ldots,X_{n})\rvert}{MAD(X_{1},\ldots,X_{n})},$
where $Med(X_{1},\ldots,X_{n})$ and $MAD(X_{1},\ldots,X_{n})$ are defined as the sample median and the median absolute deviation respectively.
$Med(X_{1},\ldots,X_{n})$ =$=\begin{cases}X_{(k)},\text{ if $n$ is odd, }k=\frac{n+1}{2}\\ \frac{X_{(k)}+X_{(k+1)}}{2},\text{ if $n$ is even, }k=\frac{n}{2}\end{cases}$
$MAD(X_{1},\ldots,X_{n})$ $=Med\left(\left\{\lvert X_{i}-Med(X_{1},\ldots,X_{n})\rvert\right\}\right).$
This induces the SDO prefiltering procedure, defined as
$F^{sdo}_{p}(S)\vcentcolon=\left\{X\in S:SDO(X,S)<p\right\}.$
The hyperparameter $p$ that can take values in the range $(0,\infty)$, with similar interpretation as in the “$p$-sigma edit” rule with different measures for location and scale.
In Figure 1(a) we calculate the learner-agnostic risk for each prefiltering procedure as a function of $\epsilon$ as follows. First, for each value of the prefiltering hyperparameter $p$ and noise parameter $m$ we calculate the worst-case distance for each Huber learner. Then, we take the minmax value by first maximizing over the noise parameter $m$ and then minimize by the prefiltering hyperparameter $p$. That way we get a guarantee on the learner-agnostic risk. Finally, we repeat this process 8 times in order to get the expected value and standard error of $R_{agn}$. For Figure 1(b) we repeat the same process but we measure the term $\max_{l\in\mathcal{L}}R_{l}-\min_{l\in\mathcal{L}}R_{l}$.
In the experiment on $\epsilon$ we have $\Delta=\left\{0.01,1.0\right\}$ and =$\mathcal{L}=\left\{H_{\delta}:\delta\in\Delta\right\}$ and we vary $\epsilon$. In the experiment on $max_{l}R_{L}(F)-min_{l}R_{l}(F)$ we fix $\epsilon=0.2$ and set $\Delta=\left\{\delta_{1},\delta_{2}\right\}$ with $\delta_{1}=0.01$ and $\delta_{2}$ varying along the x-axis. A single run of our experiments consists of generating a sample, prefiltering with a specific prefiltering procedure with a specific hyperparameter $p$ and then applying producing estimates using the downstream learners. Then we record the risk for each learner separately. Finally, we report the respective risk minimized over the prefitlering hyperparameter $p$.
B.2 Additional details for real-world data experiments
For Adult dataset, we consider seven different algorithms, as described in Section 5.1. Then, we generate two sets of fixed hyperparameters for each of them using random search, resulting in 14 different models that form the learner set. The robust version of each model is performed by training on the noisy train set and then evaluating on the noisy validation set. The objective of the respective hyperparameter search is to maximize the validation metric, either acuracy or F1 depending on noise. The brittle version of each model is performed by training two instances of the same model. For the first instance we prefilter training and validation sets together, then train on the train set and evaluate on the validation set. The second instance is trained and evaluated on untouched noisy train and validation sets respectively. The objective of the hyperparameter search is to maximize the difference between the two evaluation metrics. We follow this procedure once for each type of noise. The resulting hyperparameters are shown in Table 3 for the shortcut setting and Table 2 for the setting of label noise.
For the CIFAR-10 dataset, we use a custom CNN that we present in Table 1. All instances of our model are trained using Adam optimizer with learning rate $1e{-3}$ and batch size $128$.
For experiments in the CIFAR-10 with label noise, prefiltering is done by training the standard Resent-9 architecture using batch size 128 and Adam optimizer with learning rate $1e{-3}$ for 7 epochs. For experiments with shortcuts the model we use for prefiltering is our custom CNN, trained using batch size 128 and learning rate $1e{-3}$ for 5 epochs.
In all experiments we calculate the price of learner-agnostic prefiltering as follows. First, we conduct multiple runs where we test our prefiltering procedure with different values of the prefiltering hyperparameter $p$ on a noisy training and validation set. In each run we train all models on the prefiltered train and validation datasets and evaluate on clean test dataset. Finally, for each run we compute the best learner-specific risk over all $p$, and then we use that to calculate the price of learner-agnostic prefiltering for each of the preset prefiltering procedures. Finally, we report the smallest price of learner-agnostic prefiltering, averaged over the specific number of runs.
When we report the price of learner-agnostic prefiltering as a price as a function of subset size $k$ in Figure 2, we do so by taking all subsets of size $k$ of the initially fixed learner set and then computing for each run the average price of learner-agnostic prefiltering over all such subsets. Finally we present means and errors over all runs.
Table 1: CNN architecture (width $w=32$, #classes = 10).
Layer Operation Output Shape
Input $C\times 32\times 32$
Conv0 Conv2D($C$, $w$, $3{\times}3$, padding=1) $w\times 32\times 32$
ReLU0 ReLU $w\times 32\times 32$
Conv1 Conv2D($w$, $2w$, $3{\times}3$, padding=1) $2w\times 32\times 32$
ReLU1 ReLU $2w\times 32\times 32$
Conv2 Conv2D($2w$, $4w$, $3{\times}3$, stride=2, padding=1) $4w\times 16\times 16$
ReLU2 ReLU $4w\times 16\times 16$
Pool0 MaxPool($3{\times}3$) $4w\times 14\times 14$
Conv3 Conv2D($4w$, $4w$, $3{\times}3$, stride=2, padding=1) $4w\times 7\times 7$
ReLU3 ReLU $4w\times 7\times 7$
Pool1 AdaptiveAvgPool($1{\times}1$) $4w\times 1\times 1$
Flatten Flatten $4w$
Linear Linear($4w$, 10) 10
Model Variant Hyperparameters
xgboost robust max_depth=9, n_estimators=100, learning_rate=0.0514
adaboost robust n_estimators=50, learning_rate=0.2560
logitboost robust n_estimators=200, learning_rate=0.5875, max_depth=5
bagging robust n_estimators=10, max_samples=0.8523, max_features=0.5433
random_forest robust n_estimators=50, max_depth=None
svm robust C=38.76, kernel=rbf, gamma=auto
ffn robust hidden_size=32, epochs=15, learning_rate=0.0554, batch_size=128
xgboost brittle max_depth=3, n_estimators=200, learning_rate=0.0261
adaboost brittle n_estimators=100, learning_rate=0.2128
logitboost brittle n_estimators=50, learning_rate=0.4685, max_depth=5
bagging brittle n_estimators=10, max_samples=0.8643, max_features=0.7434
random_forest brittle n_estimators=100, max_depth=None
svm brittle C=31.35, kernel=rbf, gamma=auto
ffn brittle hidden_size=64, epochs=15, lr=0.0927, batch_size=128
Table 2: Models with respective variant and their hyperparameters. Adult dataset, label noise.
Model Variant Hyperparameters
xgb robust n_estimators=200, max_depth=3, learning_rate=0.01
adaboost robust n_estimators=100, learning_rate=1.0
logitboost robust n_estimators=50,max_depth=10
bagging robust n_estimators=10,max_samples=1.0,max_features=1.0
random_forest robust n_est=50,max_depth=5
svm robust C=9.85,kernel=rbf, gamma=scale
ffn robust learning_rate=0.0088, epochs=10,hidden_size=64, batch=32
xgb brittle n_estimators=200,max_depth=10,learning_rate=0.1
adaboost brittle n_estimators=200, learning_rate=1.0
logitboost brittle n_estimators=200,max_depth=5
bagging brittle n_estimators=100,max_samples=1.0,max_features=1.0
random_forest brittle n_estimators=200,max_depth=5
svm brittle C=10.72, kernel=rbf, gamma=scale
ffn brittle learning_rate=0.00295, epochs=10, hidden_size=64, batch=32
Table 3: Models with respective variant and their hyperparameters. Adult dataset, shortcuts.
Appendix C Additional experiments on real-world data
C.1 Dependence of $P(F)$ on learner diameter and number of models
We consider further ablation studies of the impact of learner heterogeneity on the price of learner-agnostic prefiltering. Let us parametrize each learner by the $\log_{10}$ value of its regularization parameter. We conduct the same setup as in the CIFAR-10 + shortcuts, but we now consider 25 learners whose learner parameters $c_{1},\ldots,c_{25}$ are equidistant in $\left[-4.43,-1\right]$. Then, for parameters $m$ and $d$, we fix 25 learner sets, each centered at one of the aforementioned learners, such set $i$ is centered at $c_{i}$, contains $m$ models whose parameters are equidistant in $[c_{i}-cd(m-1)/2,c_{i}+cd(m-1)/2]$, where $c=c_{2}-c_{1}=0.14$. Then we compute the average price of learner-agnostic prefiltering over all of the 25 models. By only varying one of $m$ and $d$ we can isolate the dependence on the diameter of the learners, either in terms of the parameters or the number of learner sets. Results are presented in Figure 3. We see that both the size and diameter of the learner set contribute to the increase in price of learner-agnostic prefiltering, but the contribution of the size of the learner set is more significant.
Figure 3: Price of learner-agnostic prefiltering for CIFAR-10 with shortcuts.
C.2 Learner-specific risks
In this subsection we present the ability of the prefiltering procedure to protect all learners individually. On each plot, we will present the test metric of each model in the absence of noise ($\epsilon=0$), in the presence of noise but no prefiltering($\epsilon>0,\text{no pref}$ ), and in the presence of noise and optimal learner-specific prefiltering($\epsilon>0,spec$). The results for the Adult dataset are presented as a histogram, where on x axis we have all fourteen models, each described by their algorithm and whether they are robust(r) or brittle(b). Results are shown in Figure 4 for the setting of label noise and Figure 5 for the setting of shortcuts. Note that we always use test accuracy in the settings containing label noise, and test F1 score on the settings containing shortcuts.
Figure 4: Performance of learners in learning-specific regime, Adult dataset, label noise.
Figure 5: Performance of learners in learning-specific regime, Adult dataset, shortcut.
We also plot the same metrics for the tasks on the CIFAR-10 dataset. This time, we plot all values of the regularization hyperparameter on the x-axis, and then we plot the same metrics for each learner. Results are shown in Figure 6.
(a) CIFAR-10, label noise.
(b) CIFAR-10, shortcuts.
Figure 6: Performance of perfiltering procedure in learner-specific regime, CIFAR-10 dataset.
C.3 Experiments on different prefiltering mechanisms
For the setting of CIFAR-10 with shortcuts, we also conducted a single experiment in the canonical of $\epsilon=25.5\%$, with a different prefiltering procedure. The new prefiltering procedure trains an instance of our custom CNN with early stopping. During training, the prefiltering procedure tracks the number of epochs to reach loss value $<0.01$ for each data point in the training set, and then top-$p$ fraction of the points with the lowest number of epochs to low loss value is removed. This again works under the assumption that models fit more quickly spurious correlations that stem from the shortcut patch. Results are presented in Figure 7. We also present a variation where we use a 2-layer neural network instead of our CNN. We present the results in Figure 8.
We see that the two variations of the prefiltering procedure are also effective at protecting each learner individually. Nevertheless, we observe a significant price of learner-agnostic prefiltering. We use this experiment as an indication that the significance of the price of learner-agnostic prefiltering is present across different prefiltering procedures.
(a) Price of learner-agnostic prefiltering as a function of learner heterogeneity.
(b) F1 Accuracy in the presence of no noise (blue), $\epsilon=25.5\%$ (orange), and $\epsilon=25.5\%$ with learner-specific prefiltering.
Figure 7: An instance of LARP with prefiltering procedure that computes time to low loss for our custom CNN.
(a) Price of learner-agnostic prefiltering as a function of learner heterogeneity.
(b) F1 Accuracy in the presence of no noise (blue), $\epsilon=25.5\%$ (orange), and $\epsilon=25.5\%$ with learner-specific prefiltering.
Figure 8: An instance of LARP with prefiltering procedure that computes time to low loss for a 2-layer neural network.
These plots suggest that our prefiltering mechanisms can be useful at protecting individual downstream learners. By reasoning about the usefulness in the learner-specific regime, combined with Lemma 5.1, we can argue about the general benefit of conducting learner-agnostic prefiltering in general. We consider the derivation of provable guarantees on prefiltering an exciting idea for future work.
C.4 Experiments on CIFAR-10 with shortcuts and different learner sets
In this subsection we present another variation of the canonical setup of CIFAR-10 with shortcuts. We now consider a different learning set that is again based on our custom CNN. Now, all models are parametrized by a scalar $w$ which dictates how to reweight training losses in accordance with the class imbalance present in the prefiltering set. In particular, we consider 10 models whose values for $w$ are equidistant in the range $\left[0,2\right]$, and each $w$ corresponds to a model that reweighs losses proportionally to $c_{1}^{w},\ldots,c_{10}^{w}$ where $c_{1},\ldots,c_{10}$ represent the proportions of each class present in the dataset. We present the price of learner-agnostic prefiltering as well as learner-specific risks in Figure 9.
(a) Price of learner-agnostic prefiltering as a function of learner heterogeneity.
(b) F1 Accuracy in the presence of no noise (blue), $\epsilon=25.5\%$ (orange), and $\epsilon=25.5\%$ with learner-specific prefiltering.
Figure 9: An instance of LARP with with learning set defined using loss reweighting coefficients.
Appendix D Measuring price of learner-agnostic prefiltering in the mean estimation setup
While Lemma 5.1 holds in the general setting, verifying the given condition requires knowledge of $P(F)$, which can be measured only after the prefiltering is conducted. However, in the setting of mean estimation we can use the upper bound from Theorem 4.2 as a surrogate measure for downstream performance. This allows us to provide a provable guarantee on the benefit of LARP that only uses the initial parameters of the instance.
Assume that we conduct the LARP instance presented in Section 4, with $\mathcal{D}=\mathcal{N}(\theta,\sigma^{2})$. We denote the set of Huber parameters $\Delta=\left\{\delta_{1},\ldots,\delta_{N}\right\}$ where without loss of generality $\delta_{1}\leq\delta_{2}\leq\ldots,\leq\delta_{N}$. Then we can calculate the price of learner-agnostic prefiltering explicitly:
$P(F)=\sum_{i=1}^{N}\mathcal{U}_{red}\left((\epsilon^{2}+\log(1/\delta_{0})/n)\sigma^{2}+\delta_{N}^{2},(\epsilon^{2}+\log(1/\delta_{0})/n)\sigma^{2}+\delta_{i}^{2}\right).$ (6)
We note that in this setting $P$ depends only on the learner set $\mathcal{L}$ through the set of hyperparameters $\Delta$.
Lemma D.1.
Assume that the utility reduction function $\mathcal{U}_{red}$ can be represented as
$\mathcal{U}_{red}(x,y)=\mathcal{R}(y)-\mathcal{R}(x),$
where the function $\mathcal{R}$ represents general utility from achieving particular risk. Furthermore, assume that the function $\mathcal{R}$ is $L$-Lipschitz, i.e. we have
$\lvert\mathcal{R}(x)-\mathcal{R}(y)\rvert\leq L\lvert x-y\rvert,$
for all $x,y>0$. Then, if in addition to the assumption to $n\geq\Omega\left(\log\left(1/\delta_{0}\right)\right)$, the dataset size $n$ satisfies
$n\geq\left(\frac{LN}{C(N-1)}\left(\max_{\delta\in\Delta}\delta^{2}-\min_{\delta\in\Delta}\delta^{2}\right)\right)^{1/\alpha},$
then, there is a payment scheme $(p_{i})^{i=1}_{N}$ such that, with probability $1-\delta_{0}$ over the randomness of the noisy sample, we have ${U^{(i)}_{agn}\geq U^{(i)}_{spec}}$ for all $i=1,\ldots,N$.
Proof.
Without loss of generality, assume that $\delta_{1}\leq\delta_{2}\leq\ldots\leq\delta_{N}$. This implies that $R_{u}^{(1)}\leq R_{u}^{(2)}\leq\ldots\leq R_{u}^{(N)}$. If every learner decides to do individual prefiltering, their utility will be
$U^{(i)}_{spec}=\mathcal{R}(R_{u}^{(i)})-Cn^{\alpha}.$
On the other hand, we can see that the learner-agnostic risk when all learners participate in the prefiltering and also the payments are $(p_{i})^{i=1}_{N}$ is
$U^{(i)}_{agn}=\mathcal{R}(R_{u}^{(N)})-p_{i}.$
Thus, the difference in the utilities can be calculated as
$U^{(i)}_{spec}-U^{(i)}_{agn}$ $=\mathcal{R}(R_{u}^{(i)})-Cn^{\alpha}-\mathcal{R}(R_{u}^{(N)})+p_{i}$
$\leq L\lvert R_{u}^{(i)}-R_{u}^{(N)}\rvert-Cn^{\alpha}+p_{i}$
$=L\left(\delta_{N}^{2}-\delta_{i}^{2}\right)-Cn^{\alpha}+p_{i}.$
Now, recall that $\sum_{i=1}^{N}p_{i}=Cn^{\alpha}$ by definition. Let us set
$p_{i}\propto Cn^{\alpha}-L\left(\delta^{2}_{N}-\delta^{2}_{i}\right).$
Then we have the inequality
$\sum_{i=1}^{N}Cn^{\alpha}-L\left(\delta^{2}_{N}-\delta^{2}_{i}\right)$ $\geq NCn^{\alpha}-NL\left(\delta^{2}_{N}-\delta^{2}_{1}\right)$
$\geq NCn^{\alpha}-\left(N-1\right)Cn^{\alpha}$
$\geq Cn^{\alpha},$
where the last inequality follows from our assumption that
$n\geq\left(\frac{LN}{C(N-1)}\left(\delta^{2}_{N}-\delta^{2}_{1}\right)\right)^{1/\alpha}.$
Hence we have that $p_{i}\leq Cn^{\alpha}-L\left(\delta_{N}-\delta_{i}\right)$ and hence $U^{(i)}_{spec}-U^{(i)}_{agn}\leq 0$, as desired.
Appendix E Miscellaneous
Section E.1 contains the computational requirements for the experiments. Section E.2 contains descriptions of licenses of the datasets used in the paper.
E.1 Compute resources
Conducting the experiments required around 200 GPU hours on 16 NVIDIA L4 GPUS with 24GB VRAM each for experiments on CIFAR-10, as well as around 100 CPU hours on a 64-core CPU node for experiments on mean estimation and Adults.
E.2 Licenses
CIFAR-10
The CIFAR-10 dataset is copyrighted by Alex Krizhevsky (2009) and made available free of charge under the MIT License. Source: https://www.cs.toronto.edu/~kriz/cifar.html
Adult Dataset (Census Income)
The Adult dataset is derived from the 1994 U.S. Census and is made available through the UCI Machine Learning Repository under the Creative Commons Attribution 4.0 International (CC BY 4.0) license. Source: https://archive.ics.uci.edu/ml/datasets/adult

### v3

@@
 Finally, we model the trade-off between the utility drop and the cost of repeated (learner-specific) prefiltering within a game-theoretic framework and showcase benefits of LARP for large datasets.
+Machine Learning, ICML
 1 Introduction
 The availability of large, public datasets has been crucial for the recent successes of statistical and machine learning methods.
-For example, public benchmarks such as ImageNET [13] and GLUE [59] have enabled the creation of numerous pre-trained and competitive models, readily available for adoption and fine-tuning by practitioners in any industry.
+For example, public benchmarks such as ImageNET (Deng et al., 2009) and GLUE (Wang et al., 2019) have enabled the creation of numerous pre-trained and competitive models, readily available for adoption and fine-tuning by practitioners in any industry.
 Similarly, public healthcare datasets, e.g.
 the public data on the COVID-19 pandemic released by the WHO [^1] https://data.who.int/dashboards/covid19/data, can serve as a valuable reference to medical professionals for estimating statistics about common diseases and treatments.
-Despite their attractiveness, public datasets often contain noisy, inaccurate or even maliciously manipulated data [8].
+Despite their attractiveness, public datasets often contain noisy, inaccurate or even maliciously manipulated data (Carlini et al., 2024).
 Since statistical methods are vulnerable to data contamination, such data issues can damage the accuracy of the downstream learning procedures applied on top of public data.
-For example, the Common Crawl web corpus, a widely used dataset in foundation model training, is known to contain toxic, biased, and factually incorrect content [23, 17].
-Models trained on such data, including early versions of GPT-3 [7], have consequently been shown to produce toxic or biased outputs [56].
-Another impact of public data inaccuracies manifested during the COVID-19 pandemic, as estimates of the disease mortality based on (hastily-collected) early public data on the virus spread were later reported to be underestimates [42].
+For example, the Common Crawl web corpus, a widely used dataset in foundation model training, is known to contain toxic, biased, and factually incorrect content (Gehman et al., 2020; Dodge et al., 2021).
+Models trained on such data, including early versions of GPT-3 (Brown et al., 2020), have consequently been shown to produce toxic or biased outputs (Stiennon et al., 2020).
+Another impact of public data inaccuracies manifested during the COVID-19 pandemic, as estimates of the disease mortality based on (hastily-collected) early public data on the virus spread were later reported to be underestimates (Msemburi et al., 2023).
 These issues motivate the question of whether and how public datasets should be prefiltered by the data provider, so as to protect the accuracy of downstream statistical and learning procedures (hereafter referred to as learners, for brevity) applied on top of the dataset.
-Intuitively, principled data prefiltering can improve overall model downstream quality and also aligns with recent calls for transparency in dataset creation [22] and for data-centric strategies in developing robust and ethical AI systems [39].
+Intuitively, principled data prefiltering can improve overall model downstream quality and also aligns with recent calls for transparency in dataset creation (Gebru et al., 2021) and for data-centric strategies in developing robust and ethical AI systems (Liang et al., 2022).
 However, such prefiltering poses a new technical challenge: that of designing data prefiltering algorithms that are learner-agnostic, in the sense of protecting the accuracy of a wide specification of learners simultaneously.
@@
 Our theoretical results indicate that the learner-agnostic risk generally becomes larger with increased heterogeneity of the specified learner set, leading to a loss in utility compared to the alternative of prefiltering data for each learner/use-case individually (learner-specific prefiltering).
-We term the average utility reduction across the learner set as the price of learner-agnostic prefiltering and study it via extensive experiments on the Adult [3] and CIFAR-10 [36] datasets, with label noise and spurious correlations data corruptions.
+We term the average utility reduction across the learner set as the price of learner-agnostic prefiltering and study it via extensive experiments on the Adult (Becker & Kohavi, 1996) and CIFAR-10 (Krizhevsky, 2009) datasets, with label noise and spurious correlations data corruptions.
 Finally, we model the trade-off between the price of LARP and the cost of repeated (learner-specific) prefiltering within a game-theoretic framework and showcase benefits of LARP for large datasets.
@@
 Robust statistics and learning
-The problem of learning reliable estimates/models from contaminated data has long been studied in the fields of robust statistics and machine learning [29, 11].
-Numerous works have studied various contamination models in the context of statistical estimation [28, 29, 33, 15, 32, 14] and supervised learning [33, 4, 16].
-Some commonly studied type of data corruption in machine learning include label noise [44, 48, 26, 45, 62] and spurious correlations [24, 53, 43, 51], which we consider in our experiments.
+The problem of learning reliable estimates/models from contaminated data has long been studied in the fields of robust statistics and machine learning (Huber, 2004; Cinà et al., 2023).
+Numerous works have studied various contamination models in the context of statistical estimation (Huber, 1964, 2004; Kearns & Li, 1988; Diakonikolas et al., 2019a; Kane et al., 2024; Diakonikolas & Kane, 2023) and supervised learning (Kearns & Li, 1988; Biggio et al., 2012; Diakonikolas et al., 2019b).
+Some commonly studied type of data corruption in machine learning include label noise (Natarajan et al., 2013; Patrini et al., 2017; Han et al., 2018; Northcutt et al., 2021; Zhang & Sabuncu, 2018) and spurious correlations (Geirhos et al., 2020; Shah et al., 2020; Nam et al., 2020; Sagawa et al., 2020), which we consider in our experiments.
 The main difference between our setting and robust statistics/ML is that we focus on finding robust data prefiltering procedures for protecting fixed learning algorithms, while robust ML focuses on finding suitable learning algorithms directly.
@@
 Recent advancements in data-centric machine learning emphasize the critical role of data quality in model performance.
-Surveys by [61] and [60] provide comprehensive overviews of methodologies aimed at improving data quality through validation, cleaning, and maintenance.
-One line of work studies learner-agnostic methods in the context of data valuation [31, 34].
-Another approach, which is focused on mitigating label noise, is the data pruning strategy [46].
+Surveys by (Zha et al., 2025) and (Whang et al., 2023) provide comprehensive overviews of methodologies aimed at improving data quality through validation, cleaning, and maintenance.
+One line of work studies learner-agnostic methods in the context of data valuation (Just et al., 2023; Kessler et al., 2025).
+Another approach, which is focused on mitigating label noise, is the data pruning strategy (Park et al., 2023).
 While we are also interested in prefiltering procedures that improve data quality, we study a new optimization objective for this task: that of minimizing the maximum loss of set of pre-specified learning algorithms.
 Data moderation
-A large body of literature studies methods for data quality control and moderation for ML data obtained via crowdsourcing [38, 2, 58, 54].
+A large body of literature studies methods for data quality control and moderation for ML data obtained via crowdsourcing (Lease, 2011; Awasthi et al., 2017; Vaughan, 2018; Sheng & Zhang, 2019).
 These works study specific models of data contamination, e.g.
@@
 Unlike these works, we only adopt limited assumptions on the type of corruption and seek to certify the quality of downstream learners directly.
-Another line of work studies the impact of data moderation in the context of human learning and disinformation [25, 19, 30].
+Another line of work studies the impact of data moderation in the context of human learning and disinformation (Haghtalab et al., 2021; Dwork et al., 2024; Huleihel & Refael, 2024).
 The social implications in such settings lead to many orthogonal considerations for moderators, in particular designing appropriate models of human learning, reducing polarization and diversity of the provided information.
@@
 In particular, several works have focused on interactions between data providers and model creators in ML.
-Data delegation and valuation have been studied by [9, 52, 1] with the goal of developing appropriate contract theory and fair remuneration methods for outsourcing data-related tasks.
-Other works consider strategic interactions between foundation model creators and entities performing fine-tuning [37].
+Data delegation and valuation have been studied by (Chen et al., 2022; Saig et al., 2023; Ananthakrishnan et al., 2024) with the goal of developing appropriate contract theory and fair remuneration methods for outsourcing data-related tasks.
+Other works consider strategic interactions between foundation model creators and entities performing fine-tuning (Laufer et al., 2024).
 Our work studies a concrete form of pipeline fragmentation, in which data preprocessing is performed independently of model training.
@@
 This can be further highlighted by rewriting Equation 1 as
-$R_{agn}(F)$ :$\vcentcolon=\max_{l\in\mathcal{L}}R_{l}(F)=\min_{l\in\mathcal{L}}R_{l}(F)+\left(\max_{l\in\mathcal{L}}R_{l}(F)-\min_{l\in\mathcal{L}}R_{l}(F)\right).$ (2)
+$\begin{split}R_{agn}(F)&\vcentcolon=\max_{l\in\mathcal{L}}R_{l}(F)=\\ &=\min_{l\in\mathcal{L}}R_{l}(F)+\left(\max_{l\in\mathcal{L}}R_{l}(F)-\min_{l\in\mathcal{L}}R_{l}(F)\right).\end{split}$ (2)
 The first term $\min_{l\in\mathcal{L}}R_{l}(F)$ denotes the best possible error that can be achieved by any learner in the set.
@@
 Huber contamination
-We assume the standard Huber contamination [28], in which the $\epsilon$-contaminated sample $S$ is drawn i.i.d.
+We assume the standard Huber contamination (Huber, 1964), in which the $\epsilon$-contaminated sample $S$ is drawn i.i.d.
 from a distribution $\mathcal{D}^{\prime}_{\theta}=(1-\epsilon)\mathcal{D}_{\theta}+\epsilon Q,$ where $Q\in\mathfrak{Q}$ is a noise distribution.
@@
 Huber estimators
-The set of learners $\mathcal{L}$ is a set of Huber estimators [28, 57, 50], which are parametrized by a scalar $\delta>0$.
+The set of learners $\mathcal{L}$ is a set of Huber estimators (Huber, 1964; Sun et al., 2020; Pensia et al., 2024), which are parametrized by a scalar $\delta>0$.
 They are defined as $\hat{\theta}_{\delta}(S)\vcentcolon=\operatorname*{argmin}_{\hat{\theta}\in\mathbb{R}}\sum_{x\in S}H_{\delta}(x-\hat{\theta})$, where the Huber loss $H_{\delta}$ is defined as
@@
 We now prove feasibility of LARP in the context of a scalar mean estimation for a Gaussian distribution $\mathcal{N}(\theta,\sigma^{2})$, by providing a prefiltering procedure with provable upper bounds on its learner-agnostic risk.
-This procedure is inspired by a popular informal rule for marking the tails of a sample as outliers [41].
+This procedure is inspired by a popular informal rule for marking the tails of a sample as outliers (Maronna et al., 2006).
 This can be adapted to define the following outlyingness measure:
@@
 The first term in Equation 5 is reminiscent of upper bounds in classic robust mean estimation (see e.g.
-[14]) and increases with the corruption rate $\epsilon$ and decreases with the sample size.
+(Diakonikolas & Kane, 2023)) and increases with the corruption rate $\epsilon$ and decreases with the sample size.
 The second term depends solely on the learner set and is always nonnegative.
-We note that small values of $\delta\approx 0$ yield estimators close to the sample median, which is known to enjoy guarantees similar to the first term in Equation 4 [14].
+We note that small values of $\delta\approx 0$ yield estimators close to the sample median, which is known to enjoy guarantees similar to the first term in Equation 5 (Diakonikolas & Kane, 2023).
 Therefore, if all estimators in $\mathcal{L}$ use a small value of $\delta$ (i.e.
@@
 In this subsection we empirically explore the dependence of $R_{agn}$ on $\epsilon$ and $\Delta$ with $\mathcal{D}_{\theta}=\mathcal{N}(0,1)$, by analyzing three different prefiltering procedures.
-The first one is $F_{p}^{q}$ as defined in Section 4.3, the other two are analogously defined using z-score [41] and Stahel-Donoho outlyingness (SDO) [55, 18, 63].
+The first one is $F_{p}^{q}$ as defined in Section 4.3, the other two are analogously defined using z-score (Maronna et al., 2006) and Stahel-Donoho outlyingness (SDO) (Stahel, 1981; Donoho, 1982; Zuo et al., 2004).
 Each procedure is being optimized over their own hyperparameter that controls the amount of data being removed.
@@
 In Figure 1(a) we notice that the empirical trend exhibits superlinear growth, in alignment with the $\epsilon^{2}$ term in Theorem 4.2.
-This is also consistent with lower bounds that we find in [14].
+This is also consistent with lower bounds that we find in (Diakonikolas & Kane, 2023).
 We also observe that $F^{z}$ and $F^{sdo}$ outperform $F^{q}$ in this setup.
@@
 In this section, we quantify the price of learner-agnostic prefiltering on datasets with inherent noise.
-Specifically, we conduct experiments in both the tabular and image classification settings, using two realistic noise models: i) Uniform Label Noise, which mimics scenarios where annotation is unreliable; and ii) Shortcuts [24], which simulates scenarios where training data contains spurious correlation features irrelevant to the labels, e.g.
+Specifically, we conduct experiments in both the tabular and image classification settings, using two realistic noise models: i) Uniform Label Noise, which mimics scenarios where annotation is unreliable; and ii) Shortcuts (Geirhos et al., 2020), which simulates scenarios where training data contains spurious correlation features irrelevant to the labels, e.g.
 textual information embedded in the CT image patches.
@@
 Datasets
-For image experiments we use CIFAR-10 [36], splitting its train set into train (80%) and validation (20%) sets and using the original test set.
-Tabular experiments utilize Adult [3], randomly split into train (64%), validation (16%), and test (20%) sets.
+For image experiments we use CIFAR-10 (Krizhevsky, 2009), splitting its train set into train (80%) and validation (20%) sets and using the original test set.
+Tabular experiments utilize Adult (Becker & Kohavi, 1996), randomly split into train (64%), validation (16%), and test (20%) sets.
 The validation sets are utilized to select the hyperparameters (Adult), and perform early-stopping (CIFAR-10).
@@
 In label noise experiments, we use the model accuracy as our risk metric $R$, $\epsilon=30\%$ uniform label noise on all classes, and prefilter up to $p\%$ of the train and validation sets.
-Motivated by the practical observation that complex models tend to learn useful features first before overfitting to noisy data, the prefiltered samples were selected by maximizing the loss of chosen overparametrized models (ResNet-9 [27] on CIFAR-10, and 2 layer fully connected NN on Adult) trained on the combined train and evaluation sets with early-stopping.
+Motivated by the practical observation that complex models tend to learn useful features first before overfitting to noisy data, the prefiltered samples were selected by maximizing the loss of chosen overparametrized models (ResNet-9 (He et al., 2016) on CIFAR-10, and 2 layer fully connected NN on Adult) trained on the combined train and evaluation sets with early-stopping.
 We found this procedure effective in practice, resulting in $>75\%$ of corrupted samples being filtered at the default value of $p=25\%$.
@@
 We verify that this results in $>85\%$ of prefiltered shortcut data in practice at the default $p=25\%$.
-We use the macro-F1 score [40] as our risk metric $R$ due to the class imbalance of the prefiltered datasets.
+We use the macro-F1 score (Manning, 2009) as our risk metric $R$ due to the class imbalance of the prefiltered datasets.
 Learner sets
@@
 For shortcut experiments we regularize the effect of the shortcuts on the training explicitly by reducing the gradient contribution of the shortcut patch pixels by a factor in the range $[1e{-6},1]$.
-For Adult experiments, we opt to use learners based on diverse set of machine learning model types — XGBoost [10], AdaBoost [20], LogitBoost [21], Bagging [5], RandomForest [6], SVM [12], as well as neural networks with two hidden layers and ReLU activations; in order to show that our conclusions generalize beyond families of models differing only by a single continuous parameter.
+For Adult experiments, we opt to use learners based on diverse set of machine learning model types — XGBoost (Chen & Guestrin, 2016), AdaBoost (Freund & Schapire, 1997), LogitBoost (Friedman et al., 2000), Bagging (Breiman, 1996), RandomForest (Breiman, 2001), SVM (Cortes & Vapnik, 1995), as well as neural networks with two hidden layers and ReLU activations; in order to show that our conclusions generalize beyond families of models differing only by a single continuous parameter.
 For each model type we select two sets of hyperparameters — robust and brittle, resulting in 14 total models in $\mathcal{L}$.
@@
 Our brittle models intuitively represent learners that chose their parameters for maximal accuracy, disregarding robustness completely.
-We train all neural networks using PyTorch [47] with Adam optimizer [35] and the remaining models are trained using Scikit-learn [49].
+We train all neural networks using PyTorch (Paszke et al., 2019) with Adam optimizer (Kingma & Ba, 2015) and the remaining models are trained using Scikit-learn (Pedregosa et al., 2011).
 In all shortcut experiments learners reweight their losses to account for class imbalance of the prefiltered sets.
@@
 This project was supported with computational resources provided by Google Cloud Platform (GCP).
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Supplementary Material
@@
 Proof.
-Corollary 1.15 of [14] states that if $\delta_{0}$ is at least $e^{-an}$ for sufficiently small $a$, then with probability $1-\delta_{0}$ we have
+Corollary 1.15 of (Diakonikolas & Kane, 2023) states that if $\delta_{0}$ is at least $e^{-an}$ for sufficiently small $a$, then with probability $1-\delta_{0}$ we have
 $\lvert m_{med}-\theta\rvert\leq\mathcal{O}\left(\epsilon+\sqrt{\log(1/\delta_{0})/n}\right)\sigma,$